OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that passes through (−1, 7) and (−2, −15).

OpenStudy (anonymous):

Slope is (7+15)/(-1+2) = 22/1 = 22 Then I apply point-slope formula and later reduce it to slope intercept form. y - 7 = 22 (x+1) y - 7 = 22x + 22 y = 22x + 29